A car (x) moving at a steady speed of 100kmh-1 passed the origin and a stationary car (y) the driven car (y) repeated to this by acceleration uniformly reacting a speed of 120kmh-1 in 15sec and continues at this speed until the overtake car x.
Draw av-+ graph of the motion of the two car's?
What is the distance traveled by car y of the point of overtaken car x?
1
Expert's answer
2020-11-23T10:31:24-0500
"a=\frac{120}{3.6\cdot 15}=2.22\frac{m}{s^2}"

"vt=V(t-15)+0.5V(15)\\100t=120(t-15)+0.5(120)(15)\\t=45\ s"

"s=\frac{100}{3.6}45=1250\ m"